OpenStudy (anonymous):

calc 3 help use stokes theorem to evaluate F*dR where F(x,y,z)= xyi+x^2j+z^2k and C is the intersection of paraboloid z=x^2+y^2 with plane z=y with counterclockwise orientation looking down the positive z axis.

OpenStudy (anonymous):

Ok, so you're using stoke's theorem to convert a line integral to a surface integral...

OpenStudy (anonymous):

I found curlF

OpenStudy (anonymous):

so am i still solving for x?

OpenStudy (anonymous):

Oh, I'm so stupid. I was going to evaluate the line integral. silly me.

OpenStudy (anonymous):

lol, i know you're suppose to find curl F and R(u,v)

OpenStudy (anonymous):

\[R(u,v)=u\mathbf{i}+v\mathbf{j}+\left ( u^2+v^2\right ) \mathbf{k}\]

OpenStudy (anonymous):

can you check my curlF and r(uv) curl = (0, 0, x) and R(uv) = -2u^2cosv, -2u^2sinv, u)

OpenStudy (anonymous):

and then i did the dot product between the two, and got u^2cosv after changing the X in the curl to ucosv

OpenStudy (anonymous):

Your curl is correct, but how did you do the parametrization of R?

OpenStudy (anonymous):

no clue lol, I'm not completely sure on how to do parametrization

OpenStudy (anonymous):

It looks wrong to me. If we take the x component squared plus the y component squared we get 4u^2, or 4 times the z component squared. However, we want that to equal the z component itself. Why not just use a simple parametrization like mine?

OpenStudy (anonymous):

are you differentiating the x and y components?

OpenStudy (anonymous):

Also, you need to take the surface differential \(\mathbf{R}_u \times \mathbf{R}_v\) and then you take the dot product of \(\mathbf{F} \cdot (\mathbf{R}_u \times \mathbf{R}_v)\) not \(\mathbf{F}\cdot \mathbf{R}\). I don't know why you are asking me about differentiating the x and y components, I did not do that.

OpenStudy (anonymous):

\[\mathbf{R}_u\times\mathbf{R}_v=-2u\mathbf{i}-2v\mathbf{j}+\mathbf{k}\]If you use my parametrization. Then \(\mathrm{curl}(\mathbf{F})\cdot \mathbf{N}=u\)

OpenStudy (anonymous):

but isnt the parameterization xi+yj+(x^2+y^2)k?

OpenStudy (anonymous):

oh shoot, -2u^2cosv, -2u^2sinv, u was suppose to be N

OpenStudy (anonymous):

@funkeemonk3y Yes. That is the one I used.

OpenStudy (anonymous):

k isnt the curl <0,0,x>?

OpenStudy (anonymous):

@mickifree12 OH, sorry.

OpenStudy (anonymous):

@Herp_Derp so when we change that to polar it's <0,0,ucosv>?

OpenStudy (anonymous):

my bad, i wrote the wrong thing, so is my N correct?

OpenStudy (anonymous):

so when we dot it isnt it just u^2cosv?

OpenStudy (anonymous):

idk I was just doing this all in my head...

OpenStudy (anonymous):

ok, so let's just say it was lol, time for the integration =[[

OpenStudy (anonymous):

so when you have curlF * N = <0,0,x> * <-2u^2cosv, -2u^2sinv, u> right?